6. At an interest rate of 6% annually, how long will it take for a sum to double? What if quarterly compounding or monthly compounding is applied?
Answer #1

Present value = Future value/(1+i)^n

i = interest rate per period

n= number of periods

=>

n = ln(future value/present value)/ln(1+i)

=>

a)

n = ln(2)/ln(1.06)

= 11.90 years

b)

n = ln(2)/ln(1+0.06/4)

= 46.56 quarters

= 11.64 years

c)

n = ln(2)/ln(1+0.06/12)

= 138.98 months

= 11.58 years